# No “Zero-Shot” Without Exponential Data: Pretraining Concept Frequency Determines Multimodal Model Performance

### v2

@@ 3.2 Result: Pretraining Frequency is Predictive of “Zero-Shot" Performance
We now probe the impact of concept frequency in pretraining datasets on the zero-shot performance of image-text models. We utilize the matched image-text concept frequencies for estimating frequency of concepts during pretraining. Our findings, illustrated comprehensively in Figs. 2 and 3, demonstrate the effect concept frequency has on model performance across various tasks and model types.
@@
Understanding the Plots. The plots in the main paper present text-image (CLIP) models’ zero-shot classification results using accuracy and text-to-image retrieval performance using Recall@10. Similarly, we present T2I generative models’ performance on image generation tasks using the expected aesthetics score. For the other aforementioned metrics for retrieval as well as other automated generation metrics along with human-rated scores, we find that they show similar trends, and we provide them for reference in Apps. B and C. For clarity, the data presentation is simplified from scatter plots to a cohesive line similar to work from Kandpal et al. [62] and Razeghi et al. [94]. The x-axis is log-scaled, and performance metrics are averaged within bins along this axis for ease-of-visualization of the log-linear correlation. We removed bins containing very few concepts per bin by standard IQR removal [122] following Kandpal et al. [62]. We additionally compute the pearson correlation $\rho$ for each line and provide significance results based on a two-tailed t-test [110].
@@
Key Finding: Log-linear scaling between concept frequency and zero-shot performance. Across all 16 plots, we observe a clear log-linear relationship between concept frequency and zero-shot performance. Note that these plots vary in (i) discriminative vs. generative model types, (ii) classification vs. retrieval tasks, (iii) model architecture and parameter scales, (iv) pretraining datasets with different curation methods and scales, (v) different evaluation metrics, (vi) different prompting strategies for zero-shot classification, and (vii) concept frequencies isolated only from image or text domains (additional experiments which show variation along (v) are presented in Apps. B and C, across (vi) are presented in Appx. A, and across (vii) are presented in Appx. D). The observed log-linear scaling trend persists across all seven presented dimensions. Thus, our results clearly reveal data hungry learning, i.e, a lack in current multimodal models’ ability to learn concepts from pretraining datasets in a sample-efficient manner.
4 Stress-Testing the Concept Frequency-Performance Scaling Trend
In this section, we seek to isolate the effect of concept frequency on zero-shot performance by controlling a widely known influential factor [127, 79]: similarity in distribution between pretraining and downstream test data. Additionally, we aim to validate our hypothesis further by examining the relationship between concept frequency and downstream performance on models trained on pretraining data with synthetically controlled concept distributions, images and captions.
@@ 4.1 Controlling for Similar Samples in Pretraining and Downstream Data
Figure 4: Stress-testing the log-linear scaling trends. We provide further evidence that the log-linear relationship between performance and concept frequency holds across different scenarios: (left) we control for the effect of “similarity” between the downstream test sets and pretraining datasets, and (right) we conduct experiments on an entirely synthetic pretraining distribution with no real-world text-captions or images.
@@
Motivation. Prior work has suggested that sample-level similarity between pretraining and downstream datasets impacts model performance [62, 79, 127, 94]. This leaves open the possibility that our frequency-performance results are simply an artifact of this factor, i.e., as concept frequency increases, it is likely that the pretraining dataset also contains more similar samples to the test sets. We hence investigate whether concept frequency remains predictive of downstream performance after controlling for sample-level similarity.
@@
Setup. We use the LAION-200M [10] dataset for this experiment. We first verified that a CLIP-ViT-B-32 model trained on LAION-200M dataset (used to study sample similarity in prior work [79]) exhibits a similar log-linear trend between concept frequency and zero-shot performance. Then, we use the near_pruning method from Mayilvahanan et al. [79] to eliminate 50 million samples most similar to the test sets from the pretraining LAION-200M dataset. We provide details for this in Sec. E.1. This removes the most similar samples between pretraining and test sets. We verify that this procedure influences the performance of the model drastically in performance across our aggregate classification and retrieval tasks respectively, replicating the findings of Mayilvahanan et al. [79].
@@
Key Finding: Concept Frequency still Predictive of Performance. We repeat our analysis on models trained with this controlled pretraining dataset with 150M samples, and report results on the same downstream classification and retrieval datasets in LABEL:{fig:ablation-prasanna-and-synthclip} (left). Despite the removal of the most similar samples between pretraining and test sets, we still consistently observe a clear log-linear relationship between pretraining frequency of test set concepts and zero-shot performance.
@@
Conclusion. This analysis reaffirms that, despite removing pretraining samples closely related to the test sets, the log-linear relationship between concept frequency and zero-shot performance persists. Note that this is despite substantial decreases in absolute performance, highlighting the robustness of concept frequency as a performance indicator.
Motivation. Sampling across real-world data might not result in significant differences in concept distribution, as we will later show in Sec. 5. Hence, we repeat our analysis on a synthetic dataset designed with an explicitly different concept distribution [51]. This evaluation aims to understand if pretraining concept frequency remains a significant performance predictor within a synthetic concept distribution, generalizing even on models pretrained on entirely synthetic images and captions.
Setup. The SynthCI-30M dataset [51] introduces a novel concept distribution, generating 30 million synthetic image-text pairs. Utilizing the publicly available data and models from this benchmark, we explore the relationship between concept frequency and model performance in this synthetic data regime.
Key Finding: Concept Frequency is still Predictive of Performance. We report results on models trained with their controlled dataset in LABEL:{fig:ablation-prasanna-and-synthclip} (right). We still consistently observe a clear log-linear relationship between concept frequency and zero-shot performance.
Conclusion. This consistency highlights that concept frequency is a robust indicator of model performance, extending even to entirely synthetically constructed datasets and pretraining concept distributions.
@@ 4.2 Testing Generalization to Purely Synthetic Concept and Data Distributions
Figure 5: Concept distribution of pre-training datasets is highly long-tailed. We showcase the distribution of pre-training frequencies of all concepts aggregated across all our downstream classification datasets. Across all three pre-training datasets, we observe very heavy tails. We normalize the concept frequencies and remove concepts with 0 counts for improved readability.
@@ 5 Additional Insights from Pretraining Concept Frequencies
We now present notable observations concerning the distribution of downstream concept frequencies across text, image, and text-image matched modalities in pretraining datasets.
Finding 1: Pretraining Datasets Exhibit Long-tailed Concept Distribution. Our analysis in Fig. 5 reveals an extremely long-tailed distribution of concept frequencies in pretraining datasets, with over two-thirds of concepts occurring at almost negligible frequencies relative to the size of the datasets. Our observations support the findings of past work that have noted the long-tailed distribution of large-scale language datasets [25, 88, 136]. As we observed with the log-linear trend, this distribution directly reflects disparities in performance.
Finding 2: Misalignment Between Concepts in Image-Text Pairs. We investigated the alignment of concepts within paired pretraining image-text data. Perfect image-text alignment is defined as every image-text pair containing the same concepts. Previous studies have qualitatively discussed the problem of misalignment in large image-text datasets [75, 124, 76]. Our analysis enables us to quantify this misalignment degree—for each image-text pair in the pretraining dataset, we find the concepts that are matched to the image and the text caption independently. If there are no intersecting concepts from the independent image and text hits, we count that pair as misaligned (detailed algorithm provided in Appx. G).  Tab. 4 shows the high degree of misalignment in all image-text pairs. To the best of our knowledge, this is the first attempt to explicitly quantify the degree of misalignment in pretraining image-text datasets. We release the precise misaligned image-text pairs in the pretraining datasets to enable better data curation.
Finding 3: Concept Frequencies Across Datasets are Correlated. Despite vast differences in the size (ranging from 3M to 400M samples) and curation strategies of the datasets analyzed, we discovered a surprisingly high correlation in concept frequencies across them, as presented in Tab. 4. This consistency suggests that the internet, as the common source of these datasets, naturally exhibits a long-tailed distribution, influencing any dataset derived from it to also display similar long-tailed behavior. This result inspired the “Let It Wag!” dataset.
Table 3: For each pretraining dataset, we present the number of misaligned image-text pairs and the misalignment degree: the fraction of misalignment pairs in the dataset.
[Uncaptioned image] Figure 6: Large-drops in accuracy on “Let It Wag!”. Across all 40 tested models, we note large performance drops compared to ImageNet. Further, we note that the gap in performance seems to be decreasing for higher-capacity models as demonstrated by the large positive slope (1.58) for the larger models.
@@
LAION400M 21,996,097 5.31%
Correlations CC3M CC12M YFCC15M L400M
CC3M 1.00 0.79 0.96 0.63
CC12M 1.00 0.97 0.74
YFCC15M 1.00 0.76
L400M 1.00
Table 3: For each pretraining dataset, we present the number of misaligned image-text pairs and the misalignment degree: the fraction of misalignment pairs in the dataset.
Table 4: We compute correlation in concept frequency across pretraining datasets. Despite significant differences in scale and curation, we consistently observe strong correlation.
@@ 6 Testing the Tail: Let It Wag!
Motivation. From the previous sections, we have identified a consistent long-tailed concept distribution, highlighting the scarcity of certain concepts on the web. This observation forms the basis of our hypothesis that models are likely to underperform when tested against data distributions that are heavily long-tailed. To test this, we carefully curate 290 concepts that were identified as the least frequent across all pretraining datasets. This includes concepts like an A310 aircraft, a wormsnake, and a tropical kingbird. We then use these concepts to create a classification test set, “Let It Wag!”.
Dataset Details. The “Let It Wag!” classification dataset comprises 130K test samples downloaded from the web using the method of Prabhu et al. [90]. The test samples are evenly distributed across 290 categories that represent long-tailed concepts. From the list of curated concepts, we download test set images, deduplicate them, remove outliers, and finally manually clean and hand-verify the labels.
Analysis Details. We ran both classification and image generation experiments on “Let It Wag!”. For classification, we assessed the performance of 40 text-image (CLIP) models on the “Let It Wag!” classification dataset, using an ensemble of 80 prompts from Radford et al. [91]. For the generative task, we utilized SD-XL [89] , SD-v2 [96], and Dreamlike-Photoreal-v2.0 [3] to generate images for the long-tailed concepts. For each model, we ran 50 diffusion steps, maintaining default settings for all other parameters.
@@
T2I Generation Results. We provide a qualitative analysis on image generation for assessing T2I models on rare concepts in Fig. 7. For diversity, we generate prompts using Gemini [112] (top row of generated images) and GPT4 [12] (bottom row of generated images). Green borders represent correct generations, red borders represent incorrect generations and yellow borders represent ambiguous generation. While descriptive prompting generally aids in improving the quality of generated images [52], we still observe T2I models failing to comprehend and accurately represent many concepts in our “Let It Wag!” dataset. Some failure cases involve misrepresenting activities (such as Pizza Tossing or Cricket Bowling as shown in Fig. 24), generating the wrong concept (Chuck-will’s-widow as shown in  Fig. 7 top), as well as not comprehending the concept at all (Ocarina in Fig. 7 bottom). We can see that Stable Diffusion models are prone to the long tail qualitatively—we also provide quantitative results in Sec. H.1.
@@
Figure 7: Qualitative results on the “Let It Wag!” dataset categories demonstrate failure cases of state-of-the-art T2I models on long-tailed concepts. In our experiments, we create 4 text prompts for each category using Gemini [112] and GPT4 [12] which are fed to 3 Stable Diffusion [96] models. Generation with red border is incorrect, with green border is correct and with yellow border is ambiguous. We observe that despite advances in high-fidelity image generation, there is scope for improvement for such concepts.
@@ 7 Related Work
Effect of Pre-training Data on Downstream Data. Several data-centric prior works [91, 46, 82, 42, 83, 74, 124, 125, 135, 109, 78, 92, 99, 100, 38, 26, 95] have highlighted the importance of pretraining data in affecting performance. Fang et al. [42] robustly demonstrated that pretraining data diversity is the key property underlying CLIP’s strong out-of-distribution generalisation behaviour. Similarly, Berlot-Attwell et al. [16] showed that attribute diversity is crucial for compositional generalization [60], namely systematicity [45]. Nguyen et al. [82] extended the Fang et al. [42] analysis to show that differences in data distributions can predictably change model performance, and that this behaviour can lead to effective data mixing strategies at pretraining time. Mayilvahanan et al. [79] complemented this research direction by showing that CLIP’s performance is correlated with the similarity between training and test datasets. Udandarao et al. [118] further showed that the frequency of certain visual data-types in the LAION-2B dataset was roughly correlated to the performance of CLIP models in identifying visual data-types. Our findings further pinpoint that the frequency of concept occurrences is a key indicator of performance. This complements existing research in specific areas like question-answering [62] and numerical reasoning [94] in large language models, where high train-test set similarity does not fully account for observed performance levels [127]. Concurrent to our work, Parashar et al. [86] also explore the problem of long-tailed concepts in the LAION-2B dataset and how it affects performance of CLIP models, supporting our findings. In contrast to their work, we look at count separately in image and text modalities, as well as across pretraining sets, and do a number of control experiments to thoroughly test the robustness of our result. Finally, our demonstration that the long tail yields a log-linear trend explicitly indicates exponential sample inefficiency in large-scale pretrained models.
8 Conclusions and Open Problems

### v3

@@ 3.2 Result: Pretraining Concept Frequency is Predictive of “Zero-Shot” Performance
We now probe the impact of pretraining concept frequency on “zero-shot” performance of models. Our main results, across various tasks and model types, are shown in Figs. 2 and 3.
@@
Understanding the Plots. The plots in the main paper present text-image (CLIP) models’ zero-shot classification results using accuracy and text-to-image retrieval performance using Recall@10. Similarly, we present T2I generative models’ performance on image generation tasks using the expected aesthetics score. For the other aforementioned metrics for retrieval as well as other automated generation metrics along with human-rated scores, we find that they show similar trends, and we provide them for reference in Apps. B and C. For clarity, the data presentation is simplified from scatter plots to a cohesive line, similar to work from Kandpal et al. [62] and Razeghi et al. [102]. The x-axis is log-scaled, and performance metrics are averaged within bins along this axis for ease-of-visualization of the log-linear correlation. We removed bins containing very few concepts per bin by standard IQR removal [132] following Kandpal et al. [62]. We additionally compute Pearson correlation $\rho$ [73] for each line and provide significance results based on a two-tailed t-test [118].
@@
Key Finding: Log-linear scaling between concept frequency and zero-shot performance. Across all the 16 different plots, we observe a clear log-linear relationship between pretraining concept frequency and zero-shot performance. These plots vary in (i) model type (discriminative vs. generative), (ii) task (classification vs. retrieval), (iii) model architecture and parameter count, (iv) pretraining dataset (curation methods and scales), (v) evaluation metrics, (vi) prompting strategies, and (vii) concept frequencies isolated only from image or text caption (additional experiments for (v) are presented in Apps. B and C, for (vi) are presented in Appx. A, and for (vii) are presented in Appx. D). The observed log-linear scaling trend persists across all seven presented dimensions. In some plots, we notice a slight dip at the high-frequency concepts—we analyse this in greater detail in Appx. L. Thus, taken together, our results reveal data-hungry learning, i.e, a lack in current multimodal models’ ability to learn concepts from pretraining datasets in a sample-efficient manner.
4 Stress-Testing Frequency-Performance Trends with Distributional Controls
Figure 4: Stress-testing the log-linear scaling trends. We provide further evidence for the log-linear relationship between performance and concept frequency, across different scenarios: (left) we control for “similarity” between downstream test sets and pretraining datasets, and (right) we conduct experiments on an entirely synthetic pretraining distribution with no real-world images or captions.
In this section, we perform two control experiments to account for different confounding distributional factors of pretraining datasets, to ensure the robustness of our log-linear frequency-performance scaling trends: (1) we control for sample-level similarity in distribution between pretraining and evaluation datasets [137, 81], and (2) we investigate effects of pretraining data with radically different controlled concept distributions, with entirely synthetically-generated image-text pairs [52].
@@ 4.1 Controlling for Similar Samples in Pretraining and Downstream Data
Motivation. Prior work has suggested that sample-level similarity between pretraining and downstream datasets impacts model performance [62, 81, 137, 102]. This leaves open the possibility that our frequency-performance results are simply an artifact of this factor, i.e., as concept frequency increases, it is likely that the pretraining dataset also contains more similar samples to the test sets. We hence investigate the frequency-performance trends when controlling for sample-level similarity.
@@
Setup. We use the LAION-200M [10] dataset for this experiment. We first verify that a CLIP-ViT-B-32 model pretrained on the LAION-200M dataset (used to study sample similarity in prior work [81]) exhibits a similar log-linear trend between concept frequency and zero-shot performance. Then, we use the near_pruning method from Mayilvahanan et al. [81] to eliminate 50 million samples most similar to the test sets from the pretraining LAION-200M dataset. We provide details for this in Sec. G.1. This procedure removes the most similar samples between pretraining and test sets. We verify that this procedure influences the performance of the model drastically across our aggregate classification and retrieval tasks respectively, replicating the findings of Mayilvahanan et al. [81].
@@
Key Finding: Concept Frequency is still Predictive of Performance. We repeat our analysis on models trained with this controlled pretraining dataset with 150M samples, and report results on the same downstream classification and retrieval datasets, in LABEL:{fig:ablation-prasanna-and-synthclip} (left). Despite removing the most similar samples between pretraining and test sets, we still consistently observe a clear log-linear relationship between pretraining frequency of test set concepts and zero-shot model performance.
@@
Conclusion. This analysis reaffirms that, despite removing pretraining samples closely related to the downstream evaluation datasets, the log-linear relationship between concept frequency and zero-shot performance persists. Note that this is despite substantial decreases in absolute performance, highlighting the robustness of concept frequency as a performance indicator for CLIP models.
@@
Figure 5: Concept distribution of pre-training datasets is highly long-tailed. We showcase the distribution of pretraining frequencies of all concepts aggregated across all 17 of our downstream classification datasets. Across all the pretraining datasets, we observe very heavy tails. We normalize the concept frequencies and remove concepts with 0 counts for improved readability of the plots.
Motivation. Sampling across real-world data might not result in significant differences in concept distribution, as we will later show in Sec. 5. Hence, we repeat our analysis on a synthetic dataset designed with an explicitly different concept distribution [52]. This evaluation aims to understand if pretraining concept frequency remains a significant performance predictor within a synthetic concept distribution, generalizing even for models pretrained on entirely synthetic images and text captions.
Setup. The SynthCI-30M dataset [52] introduces a novel concept distribution, generating 30 million synthetic image-text pairs. Utilizing their publicly available data and models, we explore the relationship between concept frequency and model performance in this purely synthetic data regime.
Key Finding: Concept Frequency is still Predictive of Performance. We report results for models pretrained with the controlled SynthCI-30M dataset in LABEL:{fig:ablation-prasanna-and-synthclip} (right). We still consistently observe a clear log-linear relationship between concept frequency and zero-shot model performance.
Conclusion. This consistency highlights that concept frequency is a robust indicator of model performance, extending even to entirely synthetic datasets and pretraining concept distributions.
@@ 5 Additional Insights from Pretraining Concept Frequencies
Finding 1: Pretraining Datasets Exhibit Long-tailed Concept Distributions. Our analysis in Fig. 5 reveals an extremely long-tailed distribution of concept frequencies in pretraining datasets, with over two-thirds of concepts occurring at almost negligible frequencies relative to the size of the datasets (we highlight the “head” part of this distribution in boxes). Our observations extend the findings of past work that have noted the long-tailed distribution of large-scale language datasets [14, 26, 94, 146]. As we observed with the log-linear trend, this distribution directly reflects disparities in performance.
Finding 2: Misalignment Between Concepts in Image-Text Pairs. Our concept frequency estimation pipeline enables us to investigate the alignment of concepts within paired pretraining image-text data. Perfect image-text alignment is defined as every image-text pair containing the same concepts. Previous studies have qualitatively discussed the problem of misalignment in large image-text datasets [77, 134, 78]. Our analysis enables us to quantify this misalignment degree—for each image-text pair in the pretraining dataset, we find concepts that are matched to the image and the text caption independently. If there are no intersecting concepts from the independent image and text hits, we mark that pair as misaligned (detailed algorithm shown in Appx. K).  Tab. 4 shows the high degree of misalignment in all image-text pairs (${5}{-}{36}{\%}$). To the best of our knowledge, this is the first attempt to quantify the misalignment degree in pretraining image-text datasets explicitly. We release the precise misaligned image-text samples from pretraining datasets to enable better data curation.
Finding 3: Concept Frequencies Across Datasets are Correlated. Despite vast differences in the size (3M-400M samples) and curation strategies of the pretraining datasets, we discovered a surprisingly high correlation in concept frequencies across them (see Tab. 4). This suggests that the internet, as the common source of these datasets, naturally exhibits a long-tailed distribution, influencing any dataset derived from it to display similar long-tailed behavior also. This result inspired “Let It Wag!”.
Figure 6: Large-drops in accuracy on “Let It Wag!”. Across 40 tested CLIP models, we note large performance drops compared to ImageNet. Further, the performance gap seems to decrease for high-capacity models as demonstrated by larger positive slope (1.58) for those models.
@@
LAION-400M 21,996,097 5.31%
Table 3: For each pretraining dataset, we present the number of misaligned image-text pairs and the misalignment degree: fraction of misalignment pairs.
Correlations CC-3M CC-12M YFCC-15M LAION-400M
CC-3M 1.00 0.79 0.96 0.63
CC-12M 1.00 0.97 0.74
YFCC-15M 1.00 0.76
LAION-400M 1.00
Table 4: We compute correlation in concept frequency across pretraining datasets, observing strong correlations, despite major differences in scale and curation.
@@ 6 Testing the Tail: Let It Wag!
Motivation. In the previous sections, we identified a consistent long-tailed concept distribution across pretraining datasets, highlighting the scarcity of certain concepts on the web. This observation forms the basis of our hypothesis that models likely underperform when tested against data distributions that are heavily long-tailed. To test this, we carefully curate 290 concepts identified as the least frequent across all pretraining datasets. This includes concepts like eggnog, wormsnake, and tropical kingbird. We then use these concepts to create an evaluation dataset, “Let It Wag!”.
Dataset Details. The “Let It Wag!” classification dataset comprises 130K test samples downloaded from the web using the method of Prabhu et al. [96]. The test samples are evenly distributed across 290 categories of long-tailed concepts. From the list of curated concepts, we download test set images, deduplicate them, remove outliers, and finally manually clean and hand-verify the class labels.
Analysis Details. We run both classification and image generation experiments on “Let It Wag!”. For classification, we evaluate 40 text-image (CLIP) models on the “Let It Wag!” classification dataset, using an ensemble of 80 prompts from Radford et al. [98]. For the generation task, we utilize SD-XL [95], SD-v2 [104], and Dreamlike-Photoreal-v2.0 [3], to generate images for the long-tailed concepts. For each model, we run $50$ diffusion steps, maintaining default settings for all other parameters.
@@
T2I Generation Results. We provide a qualitative analysis on image generation for assessing T2I models on the rare “Let It Wag!”concepts in Fig. 7. For enhancing image diversity, we generate prompts using Gemini [121] (top row of generated images) and GPT-4 [12] (bottom row of generated images). Green borders represent correct generations, red borders represent incorrect generations and yellow borders represent ambiguous generations. While descriptive prompting generally aids in improving the quality of generated images [53], we still observe T2I models failing to comprehend and accurately represent many concepts in our “Let It Wag!” dataset. Some failure cases involve misrepresenting activities (such as Pizza Tossing or Cricket Bowling as shown in Fig. 29), generating the wrong concept (Chuck-will’s-widow as shown in Fig. 7 (top)), as well as not comprehending the concept at all (Ocarina in Fig. 7 (bottom)). We hence show that Stable Diffusion models are prone to the long tail qualitatively—we also provide quantitative results in Sec. N.1.
@@
Figure 7: Qualitative results on “Let It Wag!” concepts demonstrate failure cases of T2I models on the long-tail. We created 4 prompts for each concept using Gemini [121] and GPT-4 [12] which are fed to 3 Stable Diffusion [104] models. Generations with red border are incorrect, green border are correct and yellow border are ambiguous. Despite advances in high-fidelity image generation, there is large scope for improvement for such long-tail concepts (quantitative results in Sec. N.1).
@@ 7 Related Work
We discuss the most relevant prior works to ours here, and defer an extended literature review to Appx. F. Past works [98, 47, 88, 43, 76, 39, 82, 81, 67] have highlighted the importance of pretraining data for improved downstream model performance. Fang et al. [43] demonstrated that pretraining data diversity is key to CLIP’s strong out-of-distribution generalisation. Nguyen et al. [88] extended this analysis to show that differences in data distributions can change model performance, enabling effective data mixing strategies for pretraining. Mayilvahanan et al. [81] complemented these works by showing that CLIP’s performance is correlated with the similarity between pretraining and test datasets. Our findings further pinpoint that the frequency of concept occurrences is a key indicator of performance. This complements existing work in areas like question-answering [62] and numerical reasoning [102] in LLMs. Concurrent to our work, Parashar et al. [92] also explore the problem of long-tailed concepts in LAION-2B and how it affects CLIP performance, supporting our findings. In contrast to their work, our demonstration that the long tail yields a log-linear trend, explicitly indicates exponential sample inefficiency in pretrained multimodal models. Additionally, contrary to their work, we index both image and text modalities, as well as span across several scales of diverse pretraining datasets. Our frequency estimation procedure on both texts and images independently, enables us to provide a more finer-grained analysis of pretraining datasets than previously studied in the literature, like (1) quantifying the misalignment between images and text captions, (2) assessing the similarity of the different pretraining data concept distributions, and (3) doing a number of control experiments to thoroughly stress-test the robustness of our log-linear scaling results.
8 Conclusion